# Do Visual-Language Maps Capture Latent Semantics?

### v1

@@ Abstract
Visual-language models (VLMs) have recently been introduced in robotic mapping by using the latent representations, i.e., embeddings, of the VLMs to represent the natural language semantics in the map. The main benefit is moving beyond a small set of human-created labels toward open-vocabulary scene understanding. While there is anecdotal evidence that maps built this way support downstream tasks, such as navigation, rigorous analysis of the quality of the maps using these embeddings is lacking. We investigate two critical properties of map quality: queryability and consistency. The evaluation of queryability addresses the ability to retrieve information from the embeddings. We investigate two aspects of consistency: intra-map consistency and inter-map consistency. Intra-map consistency captures the ability of the embeddings to represent abstract semantic classes, and inter-map consistency captures the generalization properties of the representation. In this paper, we propose a way to analyze the quality of maps created using VLMs, which forms an open-source benchmark to be used when proposing new open-vocabulary map representations. We demonstrate the benchmark by evaluating the maps created by two state-of-the-art methods, VLMaps and OpenScene, using two encoders, LSeg and OpenSeg, using real-world data from the Matterport3D data set. We find that OpenScene outperforms VLMaps with both encoders, and LSeg outperforms OpenSeg with both methods.
@@ I Introduction
Mobile robots must understand the geometry and semantics of the environment to accomplish complex tasks. The ability of maps to capture semantics has increased with the development of computer vision systems, from detecting single objects to creating pixel-wise semantic segmentation of images, yielding dense semantic maps with labels in each map cell [1]. However, most semantic segmentation methods are trained to segment the image into a small pre-selected set of categories. This imposes challenges in real-world settings when a robot operates in an environment that the categories do not fully describe [2, 3].
Visual-Language Models (VLMs) are networks that jointly train a visual and language encoder to learn a mapping from text and image inputs into a common visual-language latent space. Modern, efficient transformer architectures [4], such as CLIP [5], are trained with large data sets, which enables these models to effectively have an open vocabulary, meaning visual semantics can be matched to natural language instead of a set of selected symbols. Therefore, the exploitation of VLMs in robotic mapping could allow richer semantics to be represented and to overcome the difficulty of adapting to new environments. While this work focuses on maps that encode semantics using VLMs, mapping methods have been proposed that can incorporate other modalities as well, such as audio [6]. We will collectively refer to all these maps as latent semantic maps.
Figure 1: Maps built from visual-language models have the ability to represent complex semantics, encompassing both the class of objects and their properties. We propose a benchmark for these type of maps and evaluate the quality of different state-of-the-art methods according to the consistency of their representation and their ability to be queried.
The semantic information stored in the map must be able to be queried. While complex queries have been long possible in semantic maps by inference using object ontologies [7], semantic labels support only binary comparison as they are enumerations with no notion of distance. The promise of latent semantic maps is that queries can be evaluated by directly comparing their embeddings to the embeddings of the map with a similarity metric, yielding virtually unrestricted query vocabulary. Several successful algorithms have been built on VLMs that can perform mobile robot tasks, such as navigation [8, 9, 10]. However, the analysis from the robotic mapping perspective is left anecdotal.
In this work, we propose a way to evaluate the quality of latent semantic maps. We consider that three aspects of the map capture its quality: queryability, consistency, and the breadth of the vocabulary. We concentrate on evaluating queryability and consistency because, to the best of our knowledge, no data sets exist to evaluate robotic maps that adequately capture open vocabulary semantics. Furthermore, we evaluate the consistency of the representation, i.e., the embeddings themselves.
Understanding the possibilities and limitations of the representation is essential for bringing the methods into real-world applications. Having comparable metrics for the quality of the latent semantic maps is vital for measuring the progress in this new research field, as this allows for evaluating new methods against the state-of-the-art. For this reason, we provide an open-source benchmark [^1] The source code will be made publicly available upon acceptance of this work. against which any new open vocabulary semantic representation can be tested. To demonstrate our benchmark, we evaluate two state-of-the-art methods: VLMaps [8] and OpenScene [11], using the Matterport3D [12] data set.
@@
    We define queryability and consistency of representation as measures of mapping quality for latent semantic maps and propose metrics to quantitatively evaluate these properties;
@@
    We provide an open-source benchmark software for latent semantic maps, enabling evaluation of the progress in this new field of research;
@@
    We provide a comprehensive analysis of two state-of-the-art methods using the proposed benchmark.
@@ II Related work
II-A Maps created with visual-language embeddings
Creating maps using embeddings of VLMs, especially CLIP [5], has recently gained attention, with the proposal of methods such as VLMaps [8], OpenScene [11], NLMaps [13], and Uni-Fusion [14].
VLMaps and OpenScene create a dense grid map, where each cell is associated with a single VLM embedding. The embeddings are created from RGBD images with a language-driven semantic segmentation encoder, such as LSeg [15] or OpenSeg [16], both based on CLIP, and back-projected to the 3D map. Each cell is represented as the mean of the embeddings projected to the cell. In the original work, the 3D map in VLMaps is projected to 2D by averaging the embeddings in the $z$-axis. Additionally, OpenScene learns an encoder that directly produces CLIP embeddings from the 3D point cloud. Combining these two approaches, they propose joint 2D-3D ”ensemble” features as the final representations in the map.
Unlike VLMaps and OpenScene, NLMaps creates a feature-based map, where each feature is associated with an VLM embedding. Uni-Fusion proposes a general method to create continuous maps from any perceptual data modality, including VLM embeddings. In this paper, we focus on VLMaps and OpenScene because the NLMaps and Uni-Fusion source code were not made fully publicly available at the time of writing.
@@ II-B Open-vocabulary queries
In contrast, in this work, we evaluate queryability such that the evaluation is not restricted to semantic segmentation of the map, which imposes that the map must be partitioned. Instead, each query is evaluated separately.
II-C Analysis of maps
When assessing the quality of robotic maps, most existing works concentrate on evaluating geometry using methods ranging from simple feature descriptors [23], frequency-based methods [24], to probabilistic measures [25].
Evaluating the geometric and semantic properties jointly is only relevant when the geometry and semantics are jointly estimated, e.g., in semantic Simulatenous Localization and Mapping (SLAM). Often, the metrics are qualitative, but quantitative metrics such as the localization accuracy or map reconstruction quality compared to ground truth objects are used as well [26, 27, 28]. Like many semantic mapping works, [29, 30, 31], we focus on the evaluation of the semantic properties separately, using binary classification metrics.
@@ III Problem statement
We aim to create a benchmark to evaluate the quality of latent semantic maps. We only consider methods where the estimation of geometry and semantics are disjoint, so we concentrate on evaluating the semantic quality of the maps. We believe that three aspects of the map capture its semantic quality: queryability, consistency, and the breadth of the vocabulary. We concentrate on evaluating queryability and consistency because, to the best of our knowledge, no data sets exist to evaluate robotic maps that adequately capture open-vocabulary semantics.
The benchmark must evaluate queryability, as it is the primary way to retrieve information from the representation and, therefore, acts as the measure of two things: first, indirectly, the ability of the representation to contain relevant information, and second, the accessibility of that information. It also must address consistency, as consistent performance within and between measuring runs, times, environments, and sensors is desirable.
@@ IV-A Evaluating queryability
Because the environment consists both of background regions (e.g., floor and ceiling) and discrete object instances (e.g., a chair or a table), we propose evaluating queryability in two ways: by measuring the voxel-based coverage of the query results and the instance retrieval capability.
IV-A1 Voxel-based queryability
The first method evaluates the overall matching between the query results compared to the ground truth in a binary classification setting. This property cannot be evaluated using standard multi-class classification, as queries do not form a partition of the map. Each query produces a query result, a binary mask, over the whole map, and the same voxel might be a match of multiple queries. This is evident from the open-vocabulary perspective; e.g., a sofa might match queries ”sofa”, ”couch”, ”place to sit”, and ”soft”.
@@ IV-A1 Voxel-based queryability
    Each map $m\in\mathcal{M}$ is queried with each query $q$ in the set of queries $\mathcal{Q}$. The query result of a single query is a binary segmentation mask of the map, $\hat{\textbf{y}}_{q}=\{\hat{y}_{1},\dots,\hat{y}_{N}\}$, consisting of $N$ voxel masks $\hat{y}\in\{true,false\}$. Combined, the query results form the set of voxel-based predictions $\hat{Y}_{V}=\{\hat{\textbf{y}}_{q}\forall q\in\mathcal{Q}\}$.
@@
    For each query, the true mask $\textbf{y}_{q}$ is created from the ground truth map $m_{g}$ such, that the voxels answering query $q$ are $true$, others $false$. The true masks form a set of true labels $Y_{V}=\{\textbf{y}_{q}\forall q\in\mathcal{Q}\}$.
@@
Each comparison of a prediction $\hat{y}$ to the corresponding ground truth $y$ yields either true positive $tp=\hat{y}\land y$, true negative $tn=\neg\hat{y}\land\neg y$, false positive $fp=\hat{y}\land\neg y$, or false negative $fn=\neg\hat{y}\land y$. Using these, binary classification metrics are defined: accuracy $a=\frac{tp+tn}{tp+tn+fp+fn}$, precision $p=\frac{tp}{tp+fp}$, recall $r=\frac{tp}{tp+fn}$, F1-score $f_{1}=\frac{2tp}{2tp+fp+fn}$, and Intersection over Union (IoU) $iou=\frac{tp}{tp+fn+tp}$.
As the regions of interest are relatively small with most of the queries, true negative predictions dominate the predictions. Therefore, accuracy cannot be used as a metric, as it incorporates true negative predictions. The other metrics are not affected by true negative predictions, so we use them as the metrics for binary classification tasks in this work.
IV-A2 Instance-based queryability
The second method evaluates the capability of the map to detect and retrieve all matching objects for the query. Each instance is predicted to match or not match the query; therefore, this measures the coverage of matches within an object rather than over the whole map.
While some latent semantic maps propose a way to perform instance segmentation, not all do, so measuring instance segmentation performance is outside of the scope of this work. Instead, we assume given ground-truth instances; these can be obtained through many different approaches, using latent semantic maps alone or combining them with other maps or modalities.
@@ IV-A2 Instance-based queryability
    Given $\mathcal{M}$, $\mathcal{Q}$, and ground truth instance segmentation $\mathcal{I}$, each map $m\in\mathcal{M}$ is queried with the each query $q\in\mathcal{Q}$, each query yielding binary predicted mask $\hat{\textbf{y}}_{q}$.
@@
    For each object instance $i\in\mathcal{I}$, the corresponding voxels $\hat{\textbf{i}}_{i,q}\subset\hat{\textbf{y}}_{q}$ are selected from the predicted map.
  3. iii)
    If the majority of the voxels in $\hat{\textbf{i}}_{i,q}$ are $true$, the prediction $\hat{y}_{i,q}$ is $true$, otherwise $false$. The predictions for each instance for each query combined form the set of predictions $\hat{Y}_{I}=\{\hat{y}_{i,q}\forall i\in\mathcal{I};q\in\mathcal{Q}\}$.
  4. iv)
    Similar to the previous method, the true map $m_{q}$ is created. If the instance $i$ on the true map answers the query $q$, the true label $y_{i,q}=true$, otherwise $false$. The true labels for each instance for each query form the set of all true labels denoted $Y_{I}$.
  5. v)
    The binary classification metrics are calculated between $Y_{I}$ and $\hat{Y}_{I}$.
From $Y_{I}$ and $\hat{Y}_{I}$, the same metrics presented for the voxel-based queryability are computed for each instance.
IV-B Evaluating consistency
We further subdivide the consistency of the map into two distinct aspects, intra-map consistency and inter-map consistency, and propose a method for evaluating each.
IV-B1 Intra-map consistency
Intra-map consistency captures the similarity of embeddings across the voxels within a map, sharing semantic meaning. The hypothesis is that embeddings sharing semantic meaning are clustered together in the latent space to allow the separability of different concepts. While each object instance is in some sense unique, a consistent set of embeddings represents an abstract base class to which the instances belong.
The method consists of the following steps:
  1. i)
    Given a map $m$, and semantic label $l$, $\mathcal{E}_{l}$ is the set of embeddings corresponding to voxels in the map $m$ where the ground truth semantic label is $l$, and $\mathcal{E}_{m}$ is the set of all embeddings in the map $m$. We form a set of tuples $\mathcal{T}=\{(\mathcal{E}_{l},\mathcal{E}_{m})\forall m\in\mathcal{M};l\in\mathcal{L}\}$, where $\mathcal{L}$ is a closed-set semantic label vocabulary. For computational efficiency and to no considerable change in results, in practice, the sets $\mathcal{E}_{l}$ and $\mathcal{E}_{m}$ are subsampled. We use a subsampling ratio of $0.1$ in this work.
  2. ii)
    The average absolute deviation $d_{l}^{m}$ is calculated for $\mathcal{E}_{l}$, and $d_{m}^{m}$ for $\mathcal{E}_{m}$, for each tuple $t\in\mathcal{T}$. The average absolute deviation $d_{a}$ measures the statistical dispersion of a set, and given a set of points =$\mathcal{E}=\{e_{1},...,e_{n}\}$ it is defined as
@@ IV-B1 Intra-map consistency
    The intra-map consistency ratio $c_{l}^{m}=\frac{d_{l}^{m}}{d_{m}^{m}}$ is computed for each tuple $t\in\mathcal{T}$, with label $l$ in map $m$. This ratio represents the distinguishability of the class compared to the map average.
IV-B2 Inter-map consistency
Inter-map consistency captures the similarity of voxels with the same semantic label across different maps. The hypothesis is that embeddings within the same label are closer to each other across maps than to those with different labels with respect to a distance metric. This would imply that the objects retain their distinctiveness across maps, and therefore, the system generalizes better across different environments.
@@ IV-B2 Inter-map consistency
    For all pairs of tuples $\big{(}(\mathcal{E}_{l,1},\mathcal{E}_{m,1}),$ $(\mathcal{E}_{l,2},\mathcal{E}_{m,2})\big{)}$, $\mathcal{E}_{m,1}\neq\mathcal{E}_{m,2}$, parametric Wasserstein 2-distance $d_{w}$ is calculated between $\mathcal{E}_{l,1}$ and $\mathcal{E}_{l,2}$.
The Wasserstein $p$-distance is computationally heavy for large sets of high-dimensional embeddings. For this reason, given an n-dimensional embedding $e\in\mathbb{R}^{n}$, we approximate the set of embeddings $\mathcal{E}$ with an $n$-dimensional normal distribution $\mathcal{N}(\mu,P)$. This allows us to have a closed-form solution for the Wasserstein 2-distance
@@
TABLE I: Average results of the voxel- and instance-based queryability tests over 10 maps.
Voxel-based Instance-based
@@
OpenScene LSeg 0.623 0.617 0.630 0.457 0.258 0.252 0.264 0.149
OpenScene OpenSeg 0.580 0.575 0.587 0.412 0.261 0.255 0.267 0.150
VLMaps LSeg 0.498 0.401 0.661 0.332 0.257 0.205 0.347 0.147
VLMaps OpenSeg 0.393 0.287 0.625 0.246 0.202 0.144 0.342 0.112

### v2

# Do Visual-Language Grid Maps Capture Latent Semantics?

@@ Abstract
Visual-language models (VLMs) have recently been introduced in robotic mapping using the latent representations, i.e., embeddings, of the VLMs to represent semantics in the map. They allow moving from a limited set of human-created labels toward open-vocabulary scene understanding, which is very useful for robots when operating in complex real-world environments and interacting with humans. While there is anecdotal evidence that maps built this way support downstream tasks, such as navigation, rigorous analysis of the quality of the maps using these embeddings is missing. In this paper, we propose a way to analyze the quality of maps created using VLMs. We investigate two critical properties of map quality: queryability and distinctness. The evaluation of queryability addresses the ability to retrieve information from the embeddings. We investigate intra-map distinctness to study the ability of the embeddings to represent abstract semantic classes and inter-map distinctness to evaluate the generalization properties of the representation. We propose metrics to evaluate these properties and evaluate two state-of-the-art mapping methods, VLMaps and OpenScene, using two encoders, LSeg and OpenSeg, using real-world data from the Matterport3D data set. Our findings show that while 3D features improve queryability, they are not scale invariant, whereas image-based embeddings generalize to multiple map resolutions. This allows the image-based methods to maintain smaller map sizes, which can be crucial for using these methods in real-world deployments. Furthermore, we show that the choice of the encoder has an effect on the results. The results imply that properly thresholding open-vocabulary queries is an open problem.
@@ I Introduction
Robots must understand the geometry and semantics of the environment to accomplish complex tasks. The semantic information stored in the map must be able to be queried. While complex queries have been long possible in semantic maps by inference using object ontologies [1], semantic labels support only binary comparison as they are enumerations with no notion of distance. This imposes challenges in real-world settings when a robot operates in an environment that the categories do not fully describe [2, 3].
Figure 1: Maps built from visual-language models have the ability to represent complex semantics, encompassing both the class of objects and their properties. We propose a way to evaluate the quality of different state-of-the-art methods according to the distinctness of their semantic concepts and their ability to be queried.
Lately, grid map methods have been proposed that use Visual-Language Model (VLM) embeddings as semantics, which we will collectively refer to as latent semantic maps. Their promise is that queries can be evaluated by directly comparing the query embedding to the embeddings of the map with a similarity metric, yielding virtually unrestricted query vocabulary. Several successful algorithms have been built on VLMs that can perform mobile robot tasks, such as navigation [4, 5, 6]. However, the analysis from the robotic mapping perspective is left anecdotal.
This work aims to bridge this gap by analyzing the performance of two state-of-the-art grid mapping methods, VLMaps [4] and OpenScene [7]. We propose two metrics for this analysis: queryability and distinctness of semantic concepts. Queryability measures the ability to find matches for queries from the representation, while distinctness measures the ability to distinguish different abstract object classes. Furthermore, we explore the practical considerations of the methods by comparing changes in queryability with varying resolutions, as the map size is a function of the resolution of the map.
@@
    We define queryability and distinctness of representation as measures of mapping quality for latent semantic maps and propose metrics for quantitatively evaluating the distinctness;
@@
    We provide a comprehensive analysis of two state-of-the-art methods using the proposed metrics;
@@
    We evaluate the behavior of the maps with different resolutions;
  4. iv)
    We make our evaluation software publicly available [^1] The source code will be made publicly available upon acceptance of this work., enabling evaluation of the progress in this new field of research.
@@ II Related work
II-A Grid maps created with visual-language embeddings
Creating maps using embeddings of VLMs, especially CLIP [8], has recently gained attention, with the proposal of methods such as VLMaps [4], OpenScene [7], NLMaps [9], Uni-Fusion [10], ConceptFusion [11], ConceptGraphs [12], and HOV-SG [13].
The most common type of map used in robotics is a dense grid map. VLMaps and OpenScene have augmented these maps with open-vocabulary semantics by associating each cell with a single VLM embedding. The embeddings are created from RGBD images with a language-driven semantic segmentation encoder, such as LSeg [14] or OpenSeg [15], both based on CLIP, and back-projected to the 3D map. Each cell is represented as the mean of the embeddings projected to the cell. In the original work, the 3D map in VLMaps is projected to 2D by averaging the embeddings on the $z$-axis. Additionally, OpenScene learns an encoder that directly produces CLIP embeddings from the 3D point cloud. Combining these two approaches, OpenScene proposes joint 2D-3D “ensemble” features as the final representations in the map.
Other methods have been proposed, which do not use grid maps or do not produce embeddings on each grid cell. NLMaps creates a feature-based map, where each feature is associated with an VLM embedding. Uni-Fusion proposes a general method to create continuous maps from any perceptual data modality, including VLM embeddings. ConceptFusion builds a dense point cloud map with an embedding for each point, whereas ConceptGraphs and HOV-SG have dense point cloud maps with one embedding per object. The extension of the evaluation of representations with object-level embeddings is left for future works.
@@ II-B Open-vocabulary queries
In contrast, in this work, we evaluate queryability such that the evaluation is not restricted to semantic segmentation of the map, which imposes that the map must be partitioned. Instead, each query is evaluated separately, as it would be in real applications.
II-C Analysis of quality of maps
Most existing works assessing the quality of robotic maps concentrate on evaluating geometry using methods ranging from simple feature descriptors [22], frequency-based methods [23], to probabilistic measures [24].
Evaluating the geometric and semantic properties jointly is only relevant when the geometry and semantics are jointly estimated, e.g., in semantic Simultaneous Localization and Mapping (SLAM). Often, the metrics are qualitative, but quantitative metrics such as the localization accuracy or map reconstruction quality compared to ground truth objects are used as well [25, 26, 27]. Like many semantic mapping works, [28, 29, 30], we focus on the evaluation of the semantic properties separately, using binary classification metrics.
@@ III Problem statement
In this study, we evaluate the quality of latent semantic maps. We only consider methods where the estimation of geometry and semantics are disjoint, so we concentrate on evaluating the semantic quality of the maps, which we consider to be captured by two properties: queryability and distinctness.
Given a map $m_{i}$, an open-vocabulary query $q$, and a query function $\varphi$, a binary mask $b^{i}_{q}=\varphi(m_{i},q)$ is acquired. Given the ground truth binary mask $b^{i}_{t}$ for the same query, the queryability $Q$ with respect to $q$ and a metric $F_{q}$ is defined as $Q^{i}_{q}=F_{q}(b^{i}_{q},b^{i}_{t})$.
Given a set $\mathcal{S}^{i}_{l}$ of semantic representations of a class $l$ in map $m_{i}$, distinctness $D$ between two labels $l$ and $l^{\prime}$, from maps $m_{i}$ and $m_{i^{\prime}}$, respectively, and with respect to a metric $F_{d}$, is defined as $D^{i,\,l}_{i^{\prime},\,l^{\prime}}=F_{d}(\mathcal{S}^{i}_{l},\mathcal{S}^{i^{\prime}}_{l^{\prime}})$.
@@ IV-A Evaluating queryability
Queryability captures the ability to retrieve information from the embeddings, i.e., to find the parts of the map that answer the query. The hypothesis is that from a map with high queryability, it is possible to find accurate matches to a wide range of queries.
Queryability is evaluated by comparing the query results and ground truth in a binary classification setting. This property cannot be evaluated using standard multi-class classification, as queries do not form a partition of the map. Each query result is a binary mask over the whole map, and the same voxel might be a match of multiple queries. This is evident from the open-vocabulary perspective; e.g., a sofa might match queries “sofa”, “couch”, “place to sit”, and “soft”.
@@
    Each map $m\in\mathcal{M}$ is queried with each query $q$ in the set of queries $\mathcal{Q}$. The query result of a single query is a binary segmentation mask of the map, $\hat{\textbf{y}}_{q}=\{\hat{y}_{1},\dots,\hat{y}_{N}\}$, consisting of $N$ voxel masks $\hat{y}\in\{true,false\}$. Combined, the query results form the set of voxel-based predictions $\hat{Y}_{V}=\{\hat{\textbf{y}}_{q}|q\in\mathcal{Q}\}$.
@@
    For each query, the true mask $\textbf{y}_{q}$ is created from the ground truth map $m_{g}$ such, that the voxels answering query $q$ are $true$, others $false$. The true masks form a set of true labels $Y_{V}=\{\textbf{y}_{q}|q\in\mathcal{Q}\}$.
@@
We use precision, recall, f1-score, and Intersection over Union (IoU) as metrics. We acquire the queryability by averaging them over all of the maps and queries with respect to each metric $Q^{\mathcal{M}}_{\mathcal{Q}}=F_{q}(Y_{V},\hat{Y}_{V}),\>F_{q}\in\{p,r,f_{1},iou\}$.
With most queries, the regions of interest are relatively small, so true negatives dominate the predictions. Therefore, accuracy cannot be used as a metric, as it incorporates true negative predictions. The selected metrics are not affected by this problem.
IV-B Evaluating distinctness
We further subdivide the distinctness of the map into two distinct aspects, intra-map distinctness, and inter-map distinctness, and propose a method for evaluating each.
IV-B1 Intra-map distinctness
Intra-map distinctness captures the similarity of embeddings across the voxels within a map, sharing semantic meaning. The hypothesis is that embeddings sharing semantic meaning are clustered together in the latent space, allowing the distinguishing of different concepts.
@@ IV-B1 Intra-map distinctness
    Given the $i$-th map $m_{i}$, and semantic label $l$, $\mathcal{E}_{l}$ is the set of embeddings corresponding to voxels in the map $m_{i}$ where the ground truth semantic label is $l$, and $\mathcal{E}_{i}$ is the set of all embeddings in the map $m_{i}$. We form a set of tuples $\mathcal{T}=\{(\mathcal{E}_{l},\mathcal{E}_{i})|m_{i}\in\mathcal{M},l\in\mathcal{L}\}$, where $\mathcal{L}$ is a closed-set semantic label vocabulary. For computational efficiency and to no considerable change in results, in practice, the sets $\mathcal{E}_{l}$ and $\mathcal{E}_{i}$ are subsampled. We subsample each label separately to preserve the label distribution. We use a subsampling ratio of $0.1$ in this work.
@@
    The average absolute deviation $d_{l}^{i}$ is calculated for $\mathcal{E}_{l}$, and $d_{\mu}^{i}$ for $\mathcal{E}_{i}$, for each tuple $t\in\mathcal{T}$. The average absolute deviation $d_{a}$ measures the statistical dispersion of a set, and given a set of points =$\mathcal{E}=\{e_{1},...,e_{n}\}$ it is defined as
@@
    The intra-map distinctness ratio $c_{l}^{i}=\frac{d_{l}^{i}}{d_{\mu}^{i}}$ is computed for each tuple $t\in\mathcal{T}$, with label $l$ in map $m_{i}$. This ratio represents the distinctness $D^{i,\,l}_{i,\,\mathcal{L}}$ of the class $l$ compared to the map average.
IV-B2 Inter-map distinctness
Inter-map distinctness captures the similarity of voxels with the same semantic label across different maps. The hypothesis is that embeddings within the same label are closer to each other across maps than to those with different labels with respect to a distance metric, i.e., $D^{i,\,l}_{j,\,l}<D^{i,\,l}_{i,\,l^{\prime}}$. This would imply that the object classes retain their distinctness across maps, and therefore, the system generalizes better across different environments.
@@ IV-B2 Inter-map distinctness
    For all pairs of tuples $\big{(}(\mathcal{E}_{l_{1}},\mathcal{E}_{i_{1}}),$ $(\mathcal{E}_{l_{2}},\mathcal{E}_{i_{2}})\big{)}$, ${i_{1}}\neq{i_{2}}$, parametric Wasserstein 2-distance $d_{w}$ is calculated between $\mathcal{E}_{l_{1}}$ and $\mathcal{E}_{l_{2}}$, representing distinctness $D^{i_{1},\,l_{1}}_{i_{2},\,l_{2}}$. We subsample $\mathcal{E}_{l_{1}}$ and $\mathcal{E}_{l_{2}}$ similarly to Section IV-B1.
If the embeddings are normalized, they can be thought of as high-dimensional probability distributions. To measure the distance between the distributions of $\mathcal{E}_{l,1}$ and $\mathcal{E}_{l,2}$, we selected the Wasserstein $p$-distance, which is a common metric of the distance between probability distributions. However, it is computationally heavy for large sets of high-dimensional embeddings. For this reason, given an n-dimensional embedding $e\in\mathbb{R}^{n}$, we approximate the set of embeddings $\mathcal{E}$ with an $n$-dimensional normal distribution $\mathcal{N}(\mu,P)$. This allows us to have a closed-form solution for the Wasserstein 2-distance
@@
Figure 2: The mean map size in gigabits (Gb) as a function of the map cell size.
TABLE I: Average results of the queryability tests over 10 maps.
Semantic segmentation VLMaps querying
@@
OpenScene LSeg 0.645 0.639 0.651 0.480 0.174 0.096 0.897 0.095
OpenScene OpenSeg 0.603 0.598 0.609 0.435 0.111 0.059 0.975 0.059
VLMaps LSeg 0.541 0.529 0.553 0.375 0.425 0.297 0.757 0.271
VLMaps OpenSeg 0.495 0.526 0.469 0.333 0.367 0.256 0.700 0.226